# On the convergence of bootstrap current to the Shaing-Callen limit in stellarators

### v1

@@ 2.2 Collisionless asymptotic solutions
$g_{-1}=\sigma\eta\int\limits_{\varphi_{0}}^{\varphi_{N}}{\rm d}\varphi\;\frac{B^{2}}{B^{\varphi}}\int\limits^{\eta_{b}}_{\eta}{\rm d}\eta^{\prime}\;\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}{\rm d}\varphi\;D_{\eta}\right)^{-1}=\sigma l_{c}\int\limits^{\eta_{b}}_{\eta}{\rm d}\eta^{\prime}\;\frac{\left\langle B^{2}\right\rangle}{\left\langle|\lambda|\right\rangle}$ (45)
@@
from the trapped-passing boundary and included an arbitrary integration constant into the lower integration limit $\varphi_{b}$ being, therefore, an arbitrary constant too. This constant is the same for all classes due continuity of $g_{0}$ at class transition boundaries. In the passing region, solutions of Eqs. (19) and (23) result in
@@
and where $C_{\sigma}$ are different integration constants of Eq. (23) for $\sigma=\pm 1$.
@@
Similar to the solution of the direct problem in the passing region, the function (47) does not depend on $\varphi_{b}$. However, it has an integrable logarithmic singularity at the trapped passing boundary, $g_{0}\propto\log(\eta_{b}-\eta)$, in particular, due to such singularity of $\partial\langle|\lambda|\rangle/\partial\eta$ which, up to a constant factor, is a bounce time. Next order correction shows again that series expansion breaks down in the boundary layer, $|\eta_{b}-\eta|\leq\delta\eta\propto l_{c}^{-1/2}$.
It is convenient now to split $g_{0}$ into even and odd in $\sigma$ parts, $g_{0}=g_{0}^{\text{even}}+g_{0}^{\text{odd}}$, where $g_{0}^{\text{even}}$ is given by Eqs. (46)-(48) with replacement of $C_{\sigma}$ to a constant $C_{\text{even}}$ and $g_{0}^{\text{odd}}=\sigma C_{\text{odd}}\Theta(\eta_{b}-\eta)$ where $C_{\text{odd}}$ is another constant.
@@
For matching through the trapped-passing boundary, we take an odd part of Eq. (3) and, similar to (32), integrate it across the boundary layer ignoring the contribution of the source term and retaining only the leading order solution (45) in the collisional flux through the boundaries of integration domain,
@@
$\frac{\partial}{\partial\varphi}\int\limits_{\eta_{b}-\Delta\eta}^{\eta_{b}+\Delta\eta}{\rm d}\eta\;g_{0}^{\text{even}}(\varphi,\eta)=\left(\frac{\eta\langle B^{2}\rangle}{\langle|\lambda|\rangle}\frac{|\lambda|}{B^{\varphi}}\right)_{\eta=\eta_{b}-\Delta\eta}\rightarrow\left(\frac{\eta\langle B^{2}\rangle}{\langle|\lambda|\rangle}\frac{|\lambda|}{B^{\varphi}}\right)_{\eta=\eta_{b}},$ (49)
@@
where $\delta\eta\ll\Delta\eta\ll\eta_{b}$, $\delta\eta$ is the boundary layer width and the last expression corresponds to the limit $\Delta\eta\rightarrow 0$. Thus $g_{0}^{\text{even}}$ can be formally presented near the trapped passing boundary as
@@
$g_{0}^{\text{even}}=\delta(\eta_{b}-\eta)\int\limits_{\varphi_{b}}^{\varphi}{\rm d}\varphi^{\prime}\frac{\eta\langle B^{2}\rangle}{\langle|\lambda|\rangle}\frac{|\lambda|}{B^{\varphi}}+o(\delta\eta^{-1}),$ (50)
where the last term provides a vanishing contribution to the integral (49) and $\varphi_{b}$ is still arbitrary. To fix this constant, we notice that in case of stellarator symmetry, $B(2\varphi_{s}-\varphi)=B(\varphi)$ and $B^{\varphi}(2\varphi_{s}-\varphi)=B^{\varphi}(\varphi)$ where $\varphi_{s}$ is a symmetry point, stellarator symmetric part of $g^{\text{even}}$ is independent of $\varphi$. Namely, it can be checked that function $g^{\ast}(\varphi,\eta,\sigma)\equiv g^{\ast}(2\varphi_{s}-\varphi,\eta,-\sigma)$ satisfies the same Eq. (3) but with an opposite sign of the source $s_{(3)}$ and, therefore, $g^{\ast}=-g+C$ where $C$ is an arbitrary constant. Thus, $g^{\text{even}}(\varphi,\eta)+g^{\text{even}}(2\varphi_{s}-\varphi,\eta)=C$ which fixes $\varphi_{b}$ to $\varphi_{b}=\varphi_{s}=(\varphi_{0}+\varphi_{N})/2$ being the second symmetry point which necessarily exists besides the global maximum point $\varphi_{0}$ where the jump (49) is undetermined, and function (50) is discontinuous with $\varphi$. Since $g_{0}$ is determined up to a constant, we can add it to all three solutions (46), (47) and (50) to fix $\varphi_{b}=\varphi_{s}$ in Eq. (46). This change is negligible in the boundary layer solution (50), and it only re-defines the unknown $C_{\text{even}}$ in the passing region where we can safely set $\varphi_{b}=\varphi_{s}$ since the solution (47) does not depend on this integration limit. Thus we can combine even parts of all three solutions into
@@
$g_{0}^{\text{even}}(\varphi,\eta)=\int\limits_{\varphi_{b}}^{\varphi}{\rm d}\varphi^{\prime}\left(\frac{B^{2}}{B^{\varphi}}-\frac{\partial}{\partial\eta}\left(\Theta(\eta_{b}-\eta)\frac{\eta\langle B^{2}\rangle}{\langle|\lambda|\rangle}\frac{|\lambda|}{B^{\varphi}}\right)\right)+\Theta(\eta_{b}-\eta)\bar{g}_{0}(\eta),$ (51)
where $\bar{g}_{0}$ is given by Eq. (48) with $C_{\sigma}=C_{\text{even}}$.
In order to determine $C_{\text{even}}$, we integrate an even part of Eq. (3) over $\varphi$ in the domain $\varphi_{\text{beg}}(\eta)<\varphi<\varphi_{\text{end}}(\eta)$ where $\eta<\eta_{b}+\Delta\eta$,
$\int\limits_{\varphi_{\text{beg}}}^{\varphi_{\text{end}}}{\rm d}\varphi\sum_{\sigma=\pm 1}\frac{\partial}{\partial\eta}D_{\eta}\frac{\partial g}{\partial\eta}=\frac{\partial}{\partial\eta}\int\limits_{\varphi_{\text{beg}}}^{\varphi_{\text{end}}}{\rm d}\varphi\sum_{\sigma=\pm 1}D_{\eta}\frac{\partial g}{\partial\eta}=0.$ (52)
This equation can be integrated once with the condition of zero flux at the strongly passing boundary $\eta=0$,
$\int\limits_{\varphi_{\text{beg}}}^{\varphi_{\text{end}}}{\rm d}\varphi\frac{|\lambda|}{B^{\varphi}}\frac{\partial g^{\text{even}}}{\partial\eta}=\frac{\partial}{\partial\eta}\int\limits_{\varphi_{\text{beg}}}^{\varphi_{\text{end}}}{\rm d}\varphi\frac{|\lambda|g^{\text{even}}}{B^{\varphi}}+\frac{1}{2}\int\limits_{\varphi_{\text{beg}}}^{\varphi_{\text{end}}}{\rm d}\varphi\frac{Bg^{\text{even}}}{|\lambda|B^{\varphi}}=0.$ (53)
Integrating this condition across the boundary layer we can express the normalized number of particles in this layer,
$N_{b}=\frac{1}{2}\int\limits_{\eta_{b}-\Delta\eta}^{\eta_{b}+\Delta\eta}{\rm d}\eta\int\limits_{\varphi_{\text{beg}}}^{\varphi_{\text{end}}}{\rm d}\varphi\frac{Bg^{\text{even}}}{|\lambda|B^{\varphi}},$ (54)
via the jump in distribution function,
@@
$N_{b}=\left[\int\limits_{\varphi_{\text{beg}}}^{\varphi_{\text{end}}}{\rm d}\varphi\frac{|\lambda|g^{\text{even}}}{B^{\varphi}}\right]_{\eta_{b}+\Delta\eta}^{\eta_{b}-\Delta\eta},$ (55)
which provides a condition on $C_{\text{even}}$. Note that integrals in the r.h.s. of (55) are not divergent in the limit $\Delta\eta\rightarrow 0$, in contrast to $g_{0}^{\text{even}}$ in the passing region, $\eta=\eta_{b}-\Delta\eta$. Since stellarator symmetric part of $g^{\text{even}}$ which can only contribute to Eq. (54) is negligible in the boundary layer, $\lim\limits_{\Delta\eta\rightarrow 0}N_{b}=0$. Similarly, due to anti-symmetry of $g^{\text{even}}\approx g_{0}^{\text{even}}$ in the trapped particle domain and of the first term in Eq. (47) in the passing particle domain, the only term contributing to the condition (55) is $\bar{g}_{0}$ which is reduced in Eq. (48) to $\bar{g}_{0}=C_{\text{even}}$. Thus, condition (55) fixes $C_{\text{even}}=0$, and the final solution for the even part of $g_{0}$ is given then by Eq. (51) with $\bar{g}_{0}=0$. This solution is the same with Eq. (12) of Helander et al. (2011) except for the lower integration limit $\varphi_{b}$ which is at the second symmetry point $\varphi_{s}$ but not at the global maximum point $\varphi_{0}$.
Note that in a more general case where stellarator symmetry is absent, point $\varphi_{b}$ in Eq. (50) should satisfy
@@
$\int\limits_{\varphi_{0}}^{\varphi_{b}}\frac{{\rm d}\varphi^{\prime}}{B^{\varphi}}|\lambda|=\int\limits_{\varphi_{b}}^{\varphi_{N}}\frac{{\rm d}\varphi^{\prime}}{B^{\varphi}}|\lambda|$ (56)
in order that $N_{b}$ stays finite to fulfil the condition (55) at vanishing collisionality. In case the global maximum value of $B$ is reached at more than one point (e.g. if the field line makes few poloidal turns in a tokamak), the boundary layer splits into few independent segments limited by $\varphi_{\text{beg}}$ and $\varphi_{\text{end}}$ located at neighboring global maxima. Respectively, condition (56) with replacement $\varphi_{0}\rightarrow\varphi_{\text{beg}}$ and $\varphi_{N}\rightarrow\varphi_{\text{end}}$ must be fulfilled for each segment independently, resulting in individual values of $\varphi_{b}$ in each segment. This affects only the boundary layer and trapped particle domain since in the passing domain $g_{0}$ is independent of $\varphi_{b}$.
@@ 2.3 Finite collisionality, numerical examples
For the numerical tests, we use the drift kinetic equation solver NEO-2 (Kernbichler et al., 2008; Kasilov et al., 2014; Martitsch et al., 2014; Kernbichler et al., 2016; Kapper et al., 2016, 2018) which, generally, employs a full linearized collision operator including the relativistic effects (Kapper et al., 2018). Here, we restrict its collision operator to the Lorentz model only. The magnetic field model corresponds to a circular tokamak with concentric flux surfaces ($\beta=0$) and with a toroidal ripple-like perturbation, given in Boozer coordinates by
@@
Here, $B_{0}(r,\vartheta)$ is the unperturbed field, $B_{00}(r)$ is the $(m,n)=(0,0)$ harmonic of this field and $\varepsilon_{M}$ is the modulation amplitude. It can be checked that this field fulfills “true surface” condition (40) at all rational flux surfaces, which is a property useful for the studies of bootstrap resonances briefly discussed in Section 4.1. We fix $n=3$ and $\varepsilon_{M}=0.05$ and consider two cases of the rotational transform $\iota=1/q$ with $\iota=1/4$ and $\iota=2/5$ where the field line is closed after one and two poloidal turns, respectively.
@@
The normalized bootstrap coefficient (43) computed by NEO-2 for the above two cases as function of the normalized collisionality $\nu_{\ast}=\pi R\nu_{\perp}/v=\pi R/(2l_{c})=\pi\iota\nu^{\ast}$ where $R$ is the major radius of the magnetic axis and $\nu^{\ast}$ is the definition of Beidler et al. (2011) is compared to Shaing-Callen limit (42) in Fig. 2. It can be seen that asymptotic value is abruptly reached in both cases at rather low collisionalities $\nu_{\ast}\sim 10^{-6}$ with the collisional $\lambda_{bB}$ dropping briefly below the Shaing-Callen limit prior to convergence. Naturally, the normalized Ware pinch coefficient $\lambda_{bB}^{\dagger}$, Eq. (57), differs from bootstrap coefficient only by small numerical errors and cannot be distinguished in this plot. For the studies of the off-set of collisional bootstrap (Ware-pinch) coefficient from the asymptotic value seen at low but finite collisionality we examine the distribution function computed by NEO-2 in direct and adjoint problem separately.
@@
An odd part of the distribution function perturbation $g_{\text{odd}}$ responsible for bootstrap current in the direct problem is shown in Fig. 3.
@@
Figure 3: Odd part of the distribution function, $g_{\text{odd}}(\varphi,\eta)=(g(\varphi,\eta,1)-g(\varphi,\eta,-1))/2$, driven by $s_{(1)}$ for different normalized collisionalities $\nu_{\ast}$ (see the legend) in case $\iota=1/4$. The trapped-passing boundary is shown by a white dotted line.
@@
The distribution of parallel equilibrium current in velocity space is better seen from Fig. 4 where an integral $\int_{\eta}^{1/B}{\rm d}\eta^{\prime}\;g_{\text{odd}}$ is shown at the middle of the field line ($\varphi=\varphi_{b}=5\pi$). According to Eq. (37) value of this integral at $\eta=0$ up to a factor equals parallel equilibrium current density. It can be seen that the contribution of the trapped particle region to the parallel current is independent of collisionality despite the varying width of class transition boundary layers which carry a significant amount of this current.
@@
Figure 4: Integral $\int_{\eta}^{1/B}{\rm d}\eta^{\prime}\;g_{\text{odd}}$ as function of the lower limit (left), and its zoom near one of class transition boundaries (right) for various plasma collisionalities (see the legend) at $\varphi=5\pi$. The trapped-passing boundary is shown by a vertical dotted line.
Off-set of the current which depends on collisionality is produced in the passing particle region. Solution $g_{\text{odd}}$ in this region is determined up to arbitrary constant, which should match then the solution in the trapped-passing boundary layer.
Figure 5: Odd part of the distribution function $g_{\text{odd}}$ near the trapped-passing boundary (vertical dotted line) at $\varphi=5\pi$ for various collisionalities (see the legend).
At sufficiently low collisionalities ($\nu_{\ast}<10^{-5}$), where bootstrap current is independent of collisionality the nearest class transition boundary layer is well separated from the trapped-passing boundary (see Fig. 5), and the boundary condition for passing particles is fully determined by the last trapped orbit nearest to the boundary between trapped and passing particle domains (Boozer and Gardner, 1990). The rest of trapped particle domain has no effect on passing distribution, because “particles” generated by bounce averaged drift in the first local ripple fully annihilate in class transition boundary layer with “anti-particles” generated in the last local ripple where the sign of geodesic curvature is opposite to the sign in the first ripple. At higher collisionalities ($\nu_{\ast}\geq 10^{-5}$), where class transition boundary layer crosses trapped-passing boundary, some part of “particles” generated in the first ripple enters the co-passing domain since they cross trapped-passing boundary with positive velocity while “anti-particles” from the last ripple enter the counter-passing domain. Both create a current off-set of the same sign. Naturally, they cannot annihilate because they enter different domains.
@@
Figure 6: Even part of the distribution function, $g_{\text{even}}(\varphi,\eta)=(g(\varphi,\eta,1)+g(\varphi,\eta,-1))/2$, driven by $s_{(3)}$ for the same cases as in Fig. 3. The transition boundary between the two highest trapping classes is shown by a white dotted line.

### v3

@@ 2.2 Collisionless asymptotic solutions
$g_{-1}=\sigma\int\limits_{\varphi_{0}}^{\varphi_{N}}{\rm d}\varphi\;\frac{B^{2}}{B^{\varphi}}\int\limits^{\eta_{b}}_{\eta}{\rm d}\eta^{\prime}\;\eta^{\prime}\;\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}{\rm d}\varphi\;D_{\eta}\right)^{-1}=\sigma l_{c}\int\limits^{\eta_{b}}_{\eta}{\rm d}\eta^{\prime}\;\frac{\left\langle B^{2}\right\rangle}{\left\langle|\lambda|\right\rangle}$ (45)
@@
from the trapped-passing boundary and included an arbitrary integration constant into the lower integration limit $\varphi_{b}$ being, therefore, an arbitrary constant too. This constant is the same for all classes due to continuity of $g_{0}$ at class transition boundaries.
In the passing region, solutions of Eqs. (19) and (23) result in
@@
and where $C_{\sigma}$ are different integration constants of Eq. (23) for $\sigma=\pm 1$. Similar to the solution of the direct problem, derivative $\partial g_{0}/\partial\eta$ of the function (47) does not depend on $\varphi_{b}$ in the passing region. Actually, the lower integration limit $\varphi_{b}$ in Eqs. (47) and (48) enters $g_{0}$ via a periodic function of $\varphi_{b}$ which adds up to the integration constant $C_{\sigma}$ (this follows from explicit evaluation of $\partial g_{0}/\partial\varphi_{b}=\left(\langle B^{2}\rangle-B^{2}(\varphi_{b})\right)/B^{\varphi}(\varphi_{b})$ whose integral over the whole range $\varphi_{0}<\varphi_{b}<\varphi_{N}$ is zero). Function $g_{0}$ has an integrable logarithmic singularity at the trapped passing boundary, $g_{0}\propto\log(\eta_{b}-\eta)$, in particular, due to such singularity of $\partial\langle|\lambda|\rangle/\partial\eta$ which, up to a constant factor, is a bounce time. Next order correction shows again that series expansion breaks down in the boundary layer, $|\eta_{b}-\eta|\leq\delta\eta\propto l_{c}^{-1/2}$.
In order to fully determine $g_{0}$ in the whole phase space we must express three integration constants, $\varphi_{b}$ and $C_{\sigma}$, via a single constant which is the only degree of freedom re-defining an equilibrium Maxwellian. For that, we must match the solutions through the trapped-passing boundary where $g_{0}$ is singular, but first it is convenient to split $g_{0}$ into even and odd parts in $\sigma$, $g_{0}=g_{0}^{\text{even}}+g_{0}^{\text{odd}}$, where $g_{0}^{\text{even}}$ is given by Eqs. (46)-(48) with replacement of $C_{\sigma}$ by a constant $C_{\text{even}}$ and $g_{0}^{\text{odd}}=\sigma C_{\text{odd}}\Theta(\eta_{b}-\eta)$ where $\Theta(x)$ is a Heaviside step function, and $C_{\text{odd}}$ is another constant.
@@
Function $g_{0}^{\text{odd}}$ is of no interest in the following since it does not contribute to the Ware pinch but only provides a correction to the leading order solution (45). As for $g_{0}^{\text{even}}$, we still need to determine it formally in the boundary layer identifying there its most singular part. For this purpose, we take the odd part of Eq. (3) and, similar to (32), integrate it across the boundary layer ignoring the contribution of the source term and retaining only the leading order solution (45) in the collisional flux through the boundaries of integration domain,
@@
$\frac{\partial}{\partial\varphi}\int\limits_{\eta_{b}-\Delta\eta}^{\eta_{b}+\Delta\eta}{\rm d}\eta\;g_{0}^{\text{even}}(\varphi,\eta)=\left(\frac{\eta\langle B^{2}\rangle}{\langle|\lambda|\rangle}\frac{|\lambda|}{B^{\varphi}}\right)_{\eta=\eta_{b}-\Delta\eta}\rightarrow\left(\frac{\eta\langle B^{2}\rangle}{\langle|\lambda|\rangle}\frac{|\lambda|}{B^{\varphi}}\right)_{\eta=\eta_{b}}.$ (49)
@@
Here, $\delta\eta\ll\Delta\eta\ll\eta_{b}$, $\delta\eta$ is the boundary layer width and the last expression corresponds to the collisionless limit $\Delta\eta\rightarrow 0$. Thus $g_{0}^{\text{even}}$ can be formally presented near the trapped passing boundary as
@@
$g_{0}^{\text{even}}=\delta(\eta_{b}-\eta)\left(\int\limits_{\varphi_{b}}^{\varphi}{\rm d}\varphi^{\prime}\frac{\eta\langle B^{2}\rangle}{\langle|\lambda|\rangle}\frac{|\lambda|}{B^{\varphi}}+C_{\text{bou}}\right)+o(\delta\eta^{-1}),$ (50)
where $C_{\text{bou}}$ is yet another integration constant, and the last term provides a vanishing contribution to the $\eta$-integral in (49). Similar to the passing particle solution (47) which is defined for the whole closed field line and which is a periodic function of $\varphi$ with period $\varphi_{N}-\varphi_{0}$, boundary layer solution (50) is formally defined for the whole field period too and must be a periodic function as well. In order to be periodic, function (50) must be discontinuous over the variable $\varphi$, and the only point where a jump restoring the periodicity is possible, as we see below, is the global maximum point $\varphi=\varphi_{0}$ (and $\varphi=\varphi_{N}$). This discontinuity requirement leads to a certain paradox because according to Eq. (3) the distribution function is continuous with $\varphi$ everywhere in the range where $\eta B(\varphi)<1$, which includes the whole trapped-passing boundary approached from the passing side, $\eta=\eta_{b}-0$, what in turn seems to contradict the discontinuity of the function (50). To resolve this paradox, we note that a formal representation with a $\delta$-function actually assumes a small but finite width of the boundary layer, which is partly located in the passing particle region where the distribution is strictly continuous with $\varphi$, and partly in the trapped particle region where it is continuous too except a small forbidden region with $\eta B(\varphi)>1$ located near the global maximum. Discontinuity of $g_{0}^{\text{even}}$ within this trapped part of the boundary layer is sufficient to provide the required jumps restoring the periodicity.
Similarly to the boundary layer, the periodicity argument means for the trapped particle distribution that its integral form (46) determined via $\varphi_{b}$ from the interval $\varphi_{0}<\varphi_{b}<\varphi_{N}$ can be used only in the same interval, $\varphi_{0}<\varphi<\varphi_{N}$. Thus, an extension of Eqs. (46) and (49) to the infinite $\varphi$ range is obtained replacing there the lower integration limit $\varphi_{b}$ with $\varphi_{b}+(\varphi_{N}-\varphi_{0})\left[(\varphi-\varphi_{0})/(\varphi_{N}-\varphi_{0})\right]$ where []delimited-[][$\dots$][ … ] denotes the integer part. As already mentioned above, the passing particle solution (47) is valid for the infinite $\varphi$ range as is. We can actually combine all these periodic solutions for $g_{0}^{\text{even}}$ in one multiplying so extended trapped particle solution (46) with $\Theta(\eta-\eta_{b})$, adding an even part of the passing particle solution (47) multiplied with $\Theta(\eta_{b}-\eta)$ and then adding an extended boundary layer solution (50).
In order to eliminate from such a combined solution, all but one unknown constants out of three, $\varphi_{b}$, $C_{\text{even}}$ and $C_{\text{bou}}$, we use the stellarator symmetry of our closed field line such that $B(2\varphi_{s}-\varphi)=B(\varphi)$ and $B^{\varphi}(2\varphi_{s}-\varphi)=B^{\varphi}(\varphi)$ where $\varphi_{s}$ is a symmetry point. We use the fact that stellarator symmetric part of the generalized Spitzer function $g=g_{(3)}$ given by $g_{\text{sym}}(\varphi,\eta,\sigma)\equiv\left(g(\varphi,\eta,\sigma)+g^{\ast}(\varphi,\eta,\sigma)\right)/2$ with $g^{\ast}(\varphi,\eta,\sigma)\equiv g(2\varphi_{s}-\varphi,\eta,-\sigma)$ can only be a constant. Namely, it can be checked that $g^{\ast}$ satisfies the same Eq. (3) but with an opposite sign of the source $s_{(3)}$ and, therefore, $g^{\ast}=-g+C$ where $C$ is an arbitrary constant. Thus, $g_{\text{sym}}=C/2$ in the whole phase space. Introducing now a stellarator-antisymmetric part of the distribution function, $g_{\text{asym}}(\varphi,\eta,\sigma)\equiv\left(g(\varphi,\eta,\sigma)-g^{\ast}(\varphi,\eta,\sigma)\right)/2$, we can split our solution $g_{0}^{\text{even}}$ into stellarator-symmetric and stellarator-antisymmetric parts, $g_{0}^{\text{even}}=g_{\text{sym}}^{\text{even}}+g_{\text{asym}}^{\text{even}}$. Thus, we obtain that stellarator-antisymmetric part of the combined $g_{0}^{\text{even}}$ does not contain any unknown and is determined by
@@
$g_{\text{asym}}^{\text{even}}(\varphi,\eta)=\int\limits_{\varphi_{s}}^{\varphi}{\rm d}\varphi^{\prime}\left(\frac{B^{2}}{B^{\varphi}}-\frac{\partial}{\partial\eta}\left(\Theta(\eta_{b}-\eta)\frac{\eta\langle B^{2}\rangle}{\langle|\lambda|\rangle}\frac{|\lambda|}{B^{\varphi}}\right)\right),$ (51)
where $\varphi_{s}$ is fixed now to a second stellarator symmetry point $\varphi_{s}=(\varphi_{0}+\varphi_{N})/2$ which necessarily exists besides the global maximum point $\varphi_{0}$. Solution (51) is given here only for $\varphi_{0}<\varphi<\varphi_{N}$ and can be extended to the infinite $\varphi$ range replacing there $\varphi_{s}$ with $\varphi_{s}+(\varphi_{N}-\varphi_{0})\left[(\varphi-\varphi_{0})/(\varphi_{N}-\varphi_{0})\right]$. Such an extended expression is stellarator-antisymmetric with respect to symmetry points of both kinds (in particular, anti-symmetry with respect to field maximum points $\varphi_{0}+k(\varphi_{N}-\varphi_{0})$ where $k$ are integers is enabled by the discontinuity of the extended lower integration limit at those points) and is independent of an arbitrary constant $\varphi_{b}$.
In turn, all the unknown constants enter the stellarator-symmetric part of the combined solution $g_{0}^{\text{even}}$ which is of the form
@@
$g_{\text{sym}}^{\text{even}}$ $=$ $\int\limits_{\varphi_{b}}^{\varphi_{s}}{\rm d}\varphi^{\prime}\frac{B^{2}}{B^{\varphi}}+\Theta(\eta_{b}-\eta)\left(C_{\text{even}}-\int\limits_{\varphi_{b}}^{\varphi_{s}}{\rm d}\varphi^{\prime}\frac{\langle B^{2}\rangle}{B^{\varphi}}\right)$ (52)
@@
$+$ $\delta(\eta_{b}-\eta)\left(\int\limits_{\varphi_{b}}^{\varphi_{s}}{\rm d}\varphi^{\prime}\frac{\eta\langle B^{2}\rangle}{\langle|\lambda|\rangle}\frac{|\lambda|}{B^{\varphi}}+C_{\text{bou}}\right).$
Since this function can only be a constant, we must set $C_{\text{even}}$ and $C_{\text{bou}}$ so that they annihilate respective brackets multiplying Heaviside function and $\delta$-function which permits setting $\eta=\eta_{b}$ in the sub-integrand in the latter bracket. Finally, the remaining first term can be, as usual, absorbed into the equilibrium Maxwellian so that the whole final solution $g_{0}^{\text{even}}=g_{\text{asym}}^{\text{even}}$ is given by Eq. (51). The only difference of this solution from Eq. (12) of Helander et al. (2011) is the lower integration limit $\varphi_{s}$ which is the second symmetry point (usually global minimum) instead of a global maximum point $\varphi_{0}$.
@@ 2.3 Finite collisionality, numerical examples
For the numerical tests, we use the drift kinetic equation solver NEO-2 (Kernbichler et al., 2008; Kasilov et al., 2014; Martitsch et al., 2016; Kernbichler et al., 2016; Kapper et al., 2016, 2018) which, generally, employs a full linearized collision operator including the relativistic effects (Kapper et al., 2018). Here, we restrict its collision operator to the Lorentz model only. The magnetic field model corresponds to a circular tokamak with concentric flux surfaces ($\beta=0$) and with a toroidal ripple-like perturbation, given in Boozer coordinates by
@@
Here, $B_{0}(r,\vartheta)$ is the unperturbed field, $B_{00}(r)$ is the $(m,n)=(0,0)$ harmonic of this field and $\varepsilon_{M}$ is the modulation amplitude. It can be checked that this field fulfills “true surface” condition (40) at all rational flux surfaces, which is a property useful for the studies of bootstrap resonances briefly discussed in Section 4.1. We fix $n=3$ and $\varepsilon_{M}=0.1$ and consider two cases of the rotational transform $\iota=1/q$ with $\iota=1/4$ and $\iota=2/5$ where the field line is closed after one and two poloidal turns, respectively.
@@
The normalized bootstrap coefficient (43) computed by NEO-2 for the above two cases as function of the normalized collisionality $\nu_{\ast}=\pi R\nu_{\perp}/v=\pi R/(2l_{c})=\pi\iota\nu^{\ast}$ where $R$ is the major radius of the magnetic axis and $\nu^{\ast}$ is the definition of Beidler et al. (2011) is compared to Shaing-Callen limit (42) in Fig. 2. (The latter asymptotical limit is computed here using the code NEO (Nemov et al., 1999, 2004) which should not be confused with the drift kinetic code of Belli and Candy (Belli and Candy, 2015) having the same name.) It can be seen that the asymptotic value is abruptly reached in both cases at rather low collisionalities $\nu_{\ast}\sim 10^{-6}$ with the collisional $\lambda_{bB}$ dropping briefly below the Shaing-Callen limit prior to convergence. Naturally, the normalized Ware pinch coefficient $\lambda_{bB}^{\dagger}$, Eq. (53), differs from bootstrap coefficient only by small numerical errors and cannot be distinguished in this plot. For the studies of the off-set of collisional bootstrap (Ware-pinch) coefficient from the asymptotic value seen at low but finite collisionality we examine the distribution function computed by NEO-2 in direct and adjoint problem separately.
@@
An odd part of the distribution function perturbation $g^{\text{odd}}$ responsible for bootstrap current in the direct problem is shown in Fig. 3.
@@
Figure 3: Odd part of the distribution function, $g^{\text{odd}}(\varphi,\eta)=(g(\varphi,\eta,1)-g(\varphi,\eta,-1))/2$, driven by $s_{(1)}$ for different normalized collisionalities $\nu_{\ast}$ (see the legend) in case $\iota=1/4$. The trapped-passing boundary is shown by a white dotted line.
@@
The distribution of parallel equilibrium current in velocity space is better seen from Fig. 4 where an integral $\int_{\eta}^{1/B}{\rm d}\eta^{\prime}\;g^{\text{odd}}$ is shown at the middle of the field line ($\varphi=\varphi_{b}=5\pi$). According to Eq. (37) value of this integral at $\eta=0$ up to a factor equals parallel equilibrium current density. It can be seen that the contribution of the trapped particle region to the parallel current is independent of collisionality despite the varying width of class transition boundary layers which carry a significant amount of this current.
@@
Figure 4: Integral $\int_{\eta}^{1/B}{\rm d}\eta^{\prime}\;g^{\text{odd}}$ as function of the lower limit (left), and its zoom near one of class transition boundaries (right) for various plasma collisionalities (see the legend) at $\varphi=5\pi$. The trapped-passing boundary is shown by a vertical dotted line.
Off-set of the current which depends on collisionality is produced in the passing particle region. Solution $g^{\text{odd}}$ in this region is determined up to arbitrary constant, which should match then the solution in the trapped-passing boundary layer.
Figure 5: Odd part of the distribution function $g^{\text{odd}}$ near the trapped-passing boundary (vertical dotted line) at $\varphi=5\pi$ for various collisionalities (see the legend).
At sufficiently low collisionalities ($\nu_{\ast}<10^{-5}$), where bootstrap current is independent of collisionality the nearest class transition boundary layer is well separated from the trapped-passing boundary (see Fig. 5), and the boundary condition for passing particles is fully determined by the last trapped orbit nearest to the boundary between trapped and passing particle domains (Boozer and Gardner, 1990). The rest of trapped particle domain has no effect on passing distribution, because “particles” generated by bounce averaged drift in the first local ripple fully annihilate in class transition boundary layer with “anti-particles” generated in the last local ripple where the sign of geodesic curvature is opposite to the sign in the first ripple. At higher collisionalities ($\nu_{\ast}\geq 10^{-5}$), where the trapped-passing boundary layer crosses the class transition boundary, some part of “particles” generated in the first ripple (where source $s_{(1)}>0$) and leaving this ripple with positive velocity (in the direction of a local maximum) when changing due to collisions the trapping class, can later cross the trapped-passing boundary without mirroring and enter a co-passing particle domain. Similarly, “anti-particles” from the last ripple enter the counter-passing domain. Both create a current off-set of the same sign. Naturally, they cannot annihilate because they enter different domains.
@@
Figure 6: Even part of the distribution function, $g^{\text{even}}(\varphi,\eta)=(g(\varphi,\eta,1)+g(\varphi,\eta,-1))/2$, driven by $s_{(3)}$ for the same cases as in Fig. 3. The transition boundary between the two highest trapping classes is shown by a white dotted line.